a math 7 module has 12 pages with the following numbers of words 301,298,314,327,296,287,271,326,316,333,354 and 285.what is the standard number of words per pages


1
Expert's answer
2022-04-28T05:26:44-0400

Mean of population  = 

"\mu=\dfrac{1}{12}(301+298+314+327+296+287""+271+326+316+333+354+285)=309"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}"


"=\dfrac{1}{12}(64+121+25+324+169+484"

"+1444+289+49+576+2025+576)=\dfrac{3073}{6}"

"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{3073}{6}}\approx22.6311"